# On the Welfare of EIP-1559 with Patient Bidders

### v1

@@ 2.1 Blocks, Schedules, and Social Welfare
Given a set of transactions ${\mathcal{C}}$ (where finitely many arrive at each time step) a schedule $S$ is an assignment of transactions to times, such that a transaction cannot be assigned before it arrives). Formally, for each $i\in{\mathcal{C}}$ and $t\in\mathbb{N}$, a schedule $S=\{x_{i,t}\}_{i\in{\mathcal{C}},t\in\mathbb{N}}$ where $x_{i,t}\in\{0,1\}$ specifies, for each $i\in{\mathcal{C}},t\in\mathbb{N}$, whether transaction $i$ was scheduled at time $t$, under the constraints that for every transaction $i\in{\mathcal{C}}$:
@@
We denote the set of transaction in ${\mathcal{C}}$ with per-unit value at least $\theta$ by ${\mathcal{C}}(\theta)$. For a schedule $S=\{x_{i,t}\}_{i\in{\mathcal{C}},t\in\mathbb{N}}$ we use $Q_{t}(S,\theta)=\sum_{i\in{\mathcal{C}}(\theta)}x_{i,t}\cdot q_{i}$ to denote the total capacity of transactions with per-unit value at least $\theta$ that are schedule at time $t$. We use $Q_{[1,T]}(S,\theta)=\sum_{t=1}^{T}Q_{t}(S,\theta)$ to denote the total capacity of transactions with per-unit value at least $\theta$ that are schedule at times $1,2,\ldots,T$.
@@
[…]
@@
A schedule is said to have average block size limit $B$ with slackness $\Delta:\mathbb{N}\rightarrow\mathbb{N}$ if for any number of times steps $k$, the total size of transactions during any $k$ consecutive blocks is at most $(k+\Delta(k))\cdot B$. I.e. for any $t_{0}\leq t_{1}$ we have that $\sum_{t=t_{0}}^{t_{1}}Q_{t}(S)\leq(t_{1}-t_{0}+1+\Delta(t_{1}-t_{0}+1))\cdot B$. When $\Delta$ is a constant function (for constant $Z$ it holds that $\Delta(k)=Z$), we abuse notation and simply write $\Delta=Z$.
@@ 2.2 Online Scheduling Algorithms
The actual EIP-1559 algorithm with Ethereum’s choice of parameters is such an algorithm with target block size 15 Mega gas, maximum block size twice as large ($c=2$), price-update parameter $\eta=1/8$, $p_{1}=10^{-9}$ ETH and $p_{min}=10^{-18}$ ETH.232323 The minimum price is implicit since all price calculations are done in integer multiples of Wei ($=10^{-18}$ ETH) and the first block’s price, when moving to EIP-1559 was set to one Gwei ($=10^{-9}$ ETH). We also remark that Ethereum’s EIP-1559 algorithm approximates the price update multiplier $e^{\eta\cdot(Q_{t}-B)/B}$ by $1+{\eta\cdot(Q_{t}-B)/B}$ (using the first term of the Taylor expansion, $e^{x}\approx 1+x$). Finally, when there is more demand at the current block’s price than the maximum block size, then in Ethereum’s EIP-1559 algorithm, the transactions taken are chosen according to their offered “tip” rather than adverserially. From an algorithmic (non-strategic) point of view, as we take here, since the tips may be unrelated to the actual values of the transactions, they may indeed result in any adversarial schedule.
@@
Fix any $k\in\mathbb{N}$, and consider a sequence of $k$ consecutive blocks $[i,j]$ of length $k\geq 1$ (i.e. $j-i=k-1$). If $\sum_{t=i}^{j}Q_{t}=0$ the claim trivially holds. Otherwise, let $i^{\prime}\geq i$ be the lowest index such that $Q_{i^{\prime}}>0$. Note that $i^{\prime}\leq j$ and that $\sum_{t=i}^{j}Q_{t}=\sum_{t=i^{\prime}}^{j}Q_{t}$.
@@ 3 EIP-1559 has High Welfare for Patient Bidders
In this section we prove our main result, showing that when bidders are patient, the EIP-1559 algorithm produces schedules of near-optimal welfare, provided it is given a mild resource augmentation (that does not increase with the time horizon). As a warm up we first analyze a Greedy Online Algorithm (Section 3.1), illustrating some of the techniques we use in proving our main result. We then present our main result (Section 3.2), it proof is based on a central lemma which we prove in Section 3.3.
@@ 3.1 Warm up: a Greedy Online Algorithm
Before diving into the main theorem, we can get some intuition by proving that a simple online greedy algorithm with slightly relaxed constraints, competes with the optimum. Specifically, we show that a simple online greedy algorithm with $c=2$, slackness of $1$ and extension of $1$, has welfare at least as high as the optimum. This algorithm can be viewed as setting prices at each time, but these prices are allowed to depend on all input arriving so far. To some extent it is possible to view the EIP-1559 algorithm as attempting to mimic this online algorithm by gradually approximating its prices. This gradual approximation will take extra steps, require bounds on the per-unit valuations, and incur some losses which the simple greedy algorithm does not suffer from (as, unlike the EIP-1559 algorithm, it has full access to all parameters of the transactions that has already arrived).
@@
The proof of this lemma proceeds by splitting the algorithm’s run into two phases. Let $\alpha_{\theta}$ be the last (largest) time $t$ such that $p_{t}<\theta$ (If this never happens then we take $\alpha_{\theta}=0$). By the end of the first phase, i.e. at time $\alpha_{\theta}$, there are no pending transactions whose value per unit is at least $\theta$ (since these should have been scheduled before the transaction whose value per unit is $p_{\alpha_{\theta}}$). Thus, during the first phase, the total size of transactions with value at least $\theta$ that ALG schedules is at least as high as the total scheduled by OPT, that is, $Q_{[1,\alpha_{\theta}]}(ALG,\theta)\geq Q_{[1,\alpha_{\theta}]}(OPT,\theta)$. If $\alpha_{\theta}\geq T$, the claim follows. Otherwise, $\alpha_{\theta}<T$. Now we look at the second phase, i.e. during times steps $\alpha_{\theta}+1,\ldots,T,T+1$. For any time $t$ in this time range it holds that $p_{t}\geq\theta$, and thus every transaction scheduled by the algorithm in that range has value per unit of at least $\theta$. Since, as mentioned above, in every $Z$ consecutive blocks the algorithm schedules transactions of size at least $(Z-1)\cdot B$, we have that during the second stage (including time $T+1$) our algorithm scheduled transactions of size at least $((T+1)-\alpha_{\theta}-1)\cdot B$, while $OPT$ can schedule transactions of size at most $(T-\alpha_{\theta})\cdot B$ during the range $\alpha_{\theta}+1,\ldots,T-1,T$ (i.e without the extra step at time $(T+1)$). It follows that in the second phase, the total size of transactions with value at least $\theta$ that ALG schedules (with the extra time step) is at least as high as the total scheduled by OPT, that is, $Q_{[\alpha_{\theta}+1,T+1]}(ALG,\theta)\geq(T-\alpha_{\theta})\cdot B\geq Q_{[\alpha_{\theta}+1,T]}(OPT,\theta)$. ∎
[…]
@@ 3.2 Our Main Result
Our main result is that EIP-1559 algorithms have the attractive property that for patient bidders (with some bounds on the per-unit valuations), for any time horizon, they give close to optimal welfare with respect to the benchmark of all schedules that are restricted to maximum block size $B$, with slightly less steps.
This may be viewed as surprising since the algorithm is both restricted to online decisions, and furthermore, is restricted to be price-based, and yet competes with all offline clairvoyant schedules. On the other hand, it is not limited by a worse-case block size limit $B$, but rather block size limit $B$ holds only on the average with some slackness $\Delta$ (as shown in Proposition 2.11). Additionally, the algorithm gets an extension of additional $\Gamma$ time steps, for some appropriately chosen $\Gamma$ (that does not grow with the time $T$).
@@
1 is a corollary of a more general and stronger result that we prove. First, instead of the benchmark being only integral allocations, we can take the benchmark to be all fractional allocations (in which a transaction with per-unit value $v_{i}$ that is allocated a fraction $x_{i}\in[0,1]$ contributes $v_{i}\cdot q_{i}\cdot x_{i}$ to the value of the schedule). Second, we present tighter results for the case that $q_{max}<B$, replacing the assumption that $c>2$ with the assumption that $c>1+q_{max}/B$. Third, instead of the benchmark have worst-case block size limit $B$, we allow it to only have average block size limit $B$ with slackness $\Delta^{\prime}$. To allow for this extra slackness we will allow the algorithm an additional $\Delta^{\prime}$ time steps (effectively increasing $\Gamma$ by $\Delta^{\prime}$). Finally, we will allow the initial price $p_{1}$ to be arbitrary. When it is larger than $v_{max}$ it will imply that $\Gamma$ grows as $\ln\frac{p_{1}}{p_{min}}$.
@@
We remark that if $p_{min}>v_{min}\cdot e^{-\eta}$ then the result still hold, but with respect to schedules that are only allowed to schedule transactions with value per unit at least $p_{min}\cdot e^{\eta}$. That is, there is an additional additive loss in welfare that is bounded by the welfare of any schedule of transactions of value per unit less than $p_{min}\cdot e^{\eta}$.
The cornerstone of the proof of the main result is the next lemma which shows that for any $\theta$, the EIP-1559 algorithm schedules large enough size of transactions of per-unit value $\theta$, as long as it get a long enough extension. The theorem will be easily deduced from this bound using Lemma 2.6. For the next lemma, recall that $Q_{t}(S,\theta)$ denotes the total size of transactions with per-unit value of at least $\theta$ that were scheduled by $S$ at time $t$.
@@
The theorem easily follows the main lemma by integrating Eq. 1 over $\theta$, and then applying Lemma 2.6. By integrating Eq. 1 we conclude that
[…]
@@ 3.3 Proof of The Main Lemma
Before giving a formal proof of Lemma 3.3, let us give the intuition drawing parallels with the online algorithm above. Similarly to the previous analysis we will look at the last point in time $\alpha_{\theta}$ (to be simply called $\alpha$ below) where the price was less than $\theta$ and the quantity scheduled is at most $Q_{t}\leq c\cdot B-q_{max}$, meaning that every transaction with per-unit value at least the price was scheduled. As previously, up to that point, the EIP-1559 algorithm scheduled everything with per-unit value of at least $\theta$, so it competes with any schedule $S$ (and also with a fractional one). From this point on there are always enough pending transactions that are willing to pay at least $\theta$ per-unit (quantity larger than $c\cdot B-q_{max}$), so EIP-1559 prices will increase rapidly until they first reach $\theta$ (at a time to be called $\beta$ below) from which point they can never decrease much below $\theta$ (price will always be at least $\theta\cdot e^{-\eta}$, the minimum price when price decreases from $\theta$ in a single step). So after this point in time, the EIP-1559 algorithm fills its quota with transactions of value per unit of at least $\theta\cdot e^{-\eta}$, and since its average block size limit is the same as that of $S$, it competes with it successfully if having enough scheduling time. To compensate for the time it took for the price to rise up to $\theta$ (between $\alpha$ and $\beta$), we allow the algorithm an extended time of extra $\Gamma$ steps to schedule transactions. Additionally, we allow an extra $\Delta^{\prime}$ steps to the algorithm, in order for it to compete also with schedules that have average block size limitations with slackness $\Delta^{\prime}$. We suffer a loss of a $1-e^{-\eta}$ factor due to the possible single price decrease below $\theta$.
@@
To simplify notation we use $S_{t}(\theta)=Q_{t}(S,\theta)$ and $E_{t}(\theta)=Q_{t}(E,\theta)$. Let $Z_{t}(\theta)$ denote the total size of all transactions that arrived by time $t$ with value at least $\theta$, i.e. $i\in Z_{t}(\theta)$ if $v_{i}\geq\theta$ and $t_{i}\leq t$. It holds that
@@
Thus, in the rest of the proof we assume that demand at price $\theta$ was not necessarily met at any time $t\in[\alpha+1,T+\Gamma]$, and move to lower bound the total capacity scheduled by the EIP-1559 algorithm between time $\alpha+1$ and $T+\Gamma$. As demand at price $\theta$ was not necessarily met at any time $t\in[\alpha+1,T+\Gamma]$, then for $t$ in this range of times, whenever $p_{t}\leq\theta$ then $Q_{t}>c^{\prime}\cdot B\geq B$. We conclude that if $p_{t}\leq\theta$ then $p_{t+1}=p_{t}\cdot e^{\eta\cdot(Q_{t}-B)/B}>p_{t}\cdot e^{\eta\cdot(c^{\prime}-1)}$. Thus, if $p_{\alpha}<\theta$ then price will start rising till at some time $t\geq\alpha+1$ it will holds that $p_{t}\geq\theta\geq\theta^{\prime}$.
@@
Now, we use $\beta\in[\alpha+1,T+\Gamma]$ to denote the first time the price was at least $\theta^{\prime}$ (the existence of such a time is implied by the arguments below since $\Gamma$ is large enough). So, if $\beta>1$ then it holds that $p_{\beta}\geq\theta^{\prime}$ and $p_{\beta-1}<\theta^{\prime}$ (note that for any time $t\in[\beta,T+\Gamma]$ it holds that $p_{t}\geq\theta^{\prime}$).
@@
Observe that the leftmost inequality is trivial as $\beta\geq\alpha+1$. Let us now consider the rightmost inequality. It holds due to the average block size limitation constraint of $B$ with slackness $\Delta^{\prime}$, which state that $\sum_{t=\alpha+1}^{T}S_{t}(\theta)\leq(T-\alpha+\Delta^{\prime})\cdot B$. The equality $\sum_{t=\beta}^{T+\Gamma}E_{t}(\theta^{\prime})=\sum_{t=\beta}^{T+\Gamma}Q_{t}$ holds as for any time $t\in[\beta,T+\Gamma]$ it holds that $p_{t}\geq\theta^{\prime}$, and thus all capacity is used to schedule transaction with value per unit at least $\theta^{\prime}$.
@@
Let us first consider the case that $\beta=1$ and $\alpha=0$. In this case $p_{\beta}=p_{1}$ and $p_{{T+\Gamma+1}}\geq p_{min}$. So
@@
where the left inequality holds by the definition of $\Gamma$, and the claim follows. Thus, in the rest of the proof we assume that $\beta>1$.
Recall that $\beta\in[\alpha+1,T+\Gamma]$ is the first time such $p_{\beta}\geq\theta^{\prime}$. Thus, as $\beta>1$ it holds that $p_{\beta-1}<\theta^{\prime}$ (note that we have shown that for any time $t\in[\beta,T+\Gamma]$ it holds that $p_{t}\geq\theta^{\prime}$). Thus for $\beta>1$ it holds that $p_{\beta-1}<\theta^{\prime}$, and as $p_{\alpha+1}\geq p_{min}$ we get
@@
As $\eta>0$ and $c>c^{\prime}>1$ (so $c-2>-1)$ it holds that $v_{max}\cdot e^{\eta\cdot(c-2)}>v_{max}\cdot e^{-\eta}\geq\theta\cdot e^{-\eta}\geq\theta^{\prime}$, by Eq. 4

### v2

@@ 2.1 Blocks, Schedules, and Social Welfare
Given a set of transactions $\mathcal{C}$} (where finitely many arrive at each time step) a schedule $S$ is an assignment of transactions to times, such that a transaction cannot be assigned before it arrives. Formally, for each $i\in{\mathcal{C}}$ and $t\in\mathbb{N}$, a schedule $S=\{x_{i,t}\}_{i\in{\mathcal{C}},t\in\mathbb{N}}$ where $x_{i,t}\in\{0,1\}$ specifies, for each $i\in{\mathcal{C}},t\in\mathbb{N}$, whether transaction $i$ was scheduled at time $t$, under the constraints that for every transaction $i\in{\mathcal{C}}$:
@@
We denote the set of transactions in $\mathcal{C}$} with per-unit value at least $\theta$ by ${\mathcal{C}}(\theta)$. For a schedule $S=\{x_{i,t}\}_{i\in{\mathcal{C}},t\in\mathbb{N}}$ we use $Q_{t}(S,\theta)=\sum_{i\in{\mathcal{C}}(\theta)}x_{i,t}\cdot q_{i}$ to denote the total capacity of transactions with per-unit value at least $\theta$ that are scheduled at time $t$. We use $Q_{[1,T]}(S,\theta)=\sum_{t=1}^{T}Q_{t}(S,\theta)$ to denote the total capacity of transactions with per-unit value at least $\theta$ that are scheduled at times $1,2,\ldots,T$.
@@
By definition $Q_{[1,T]}(S,\theta)=\sum_{i\in{\mathcal{C}}(\theta),t\in[T]}x_{i,t}\cdot q_{i}=\sum_{i\in{\mathcal{C}},t\in[T]}x_{i,t}\cdot q_{i}\cdot\mathbbm{1}_{v_{i}\geq\theta}$ (notice that the summation over $i$ in the first summation is only over transactions with per-unit value of at least $\theta$, while in the second summation it is over all transactions). Integrating by $\theta$ we have $\int_{0}^{\infty}Q_{[1,T]}(S,\theta)d\theta=\sum_{i\in{\mathcal{C}},t\in[T]}x_{i,t}\cdot q_{i}\cdot\int_{0}^{\infty}\mathbbm{1}_{v_{i}\geq\theta}(\theta)d\theta=\sum_{i\in{\mathcal{C}},t\in[T]}x_{i,t}\cdot q_{i}\cdot v_{i}$. Now notice that this is exactly the definition of $SW(S,[1,T])$. ∎
@@
A schedule is said to have average block size limit $B$ with slackness $\Delta:\mathbb{N}\rightarrow\mathbb{N}$ if for any number of time steps $k$, the total size of transactions during any $k$ consecutive blocks is at most $(k+\Delta(k))\cdot B$. I.e. for any $t_{0}\leq t_{1}$ we have that $\sum_{t=t_{0}}^{t_{1}}Q_{t}(S)\leq(t_{1}-t_{0}+1+\Delta(t_{1}-t_{0}+1))\cdot B$. When $\Delta$ is a constant function (for constant $Z$ it holds that $\Delta(k)=Z$), we abuse notation and simply write $\Delta=Z$.
@@ 2.2 Online Scheduling Algorithms
The actual EIP-1559 algorithm with Ethereum’s choice of parameters is such an algorithm with target block size 18 Mega gas, maximum block size twice as large ($c=2$), price-update parameter $\eta=1/8$, $p_{1}=10^{-9}$ ETH and $p_{min}=10^{-18}$ ETH. [^23] The minimum price is implicit since all price calculations are done in integer multiples of Wei ($10^{-18}$ ETH) and the first block’s price, when moving to EIP-1559 was set to one Gwei ($10^{-9}$ ETH). We also remark that Ethereum’s EIP-1559 algorithm approximates the price update multiplier $e^{\eta\cdot(Q_{t}-B)/B}$ by $1+{\eta\cdot(Q_{t}-B)/B}$ (using the first term of the Taylor expansion, $e^{x}\approx 1+x$). Finally, when there is more demand at the current block’s price than the maximum block size, then in Ethereum’s EIP-1559 algorithm, the transactions taken are chosen according to their offered “tip” rather than adversarially. From an algorithmic (non-strategic) point of view, as we take here, since the tips may be unrelated to the actual values of the transactions, they may indeed result in any adversarial schedule.
@@
Fix any $k\in\mathbb{N}$, and consider a sequence of $k$ consecutive blocks $[i,j]$ of length $k\geq 1$ (i.e. ji=k1j-i=k-1). If $\sum_{t=i}^{j}Q_{t}=0$ the claim trivially holds. Otherwise, let $i^{\prime}\geq i$ be the lowest index such that $Q_{i^{\prime}}>0$. Note that $i^{\prime}\leq j$ and that $\sum_{t=i}^{j}Q_{t}=\sum_{t=i^{\prime}}^{j}Q_{t}$.
@@ 3 EIP-1559 has High Welfare for Patient Bidders
In this section we prove our main result, showing that when bidders are patient, the EIP-1559 algorithm produces schedules of near-optimal welfare, provided it is given a mild resource augmentation (that does not increase with the time horizon). As a warm up we first analyze a Greedy Online Algorithm (Section 3.1), illustrating some of the techniques we use in proving our main result. We then present our main result (Section 3.2); its proof is based on a central lemma which we prove in Section 3.3.
@@ 3.1 Warm up: a Greedy Online Algorithm
Before diving into the main theorem, we can get some intuition by proving that a simple online greedy algorithm with slightly relaxed constraints, competes with the optimum. Specifically, we show that a simple online greedy algorithm with $c=2$, slackness of $1$ and extension of $1$, has welfare at least as high as the optimum. This algorithm can be viewed as setting prices at each time, but these prices are allowed to depend on all input arriving so far. To some extent it is possible to view the EIP-1559 algorithm as attempting to mimic this online algorithm by gradually approximating its prices. This gradual approximation will take extra steps, require bounds on the per-unit valuations, and incur some losses which the simple greedy algorithm does not suffer from (as, unlike the EIP-1559 algorithm, it has full access to all parameters of the transactions that have already arrived).
@@
The proof of this lemma proceeds by splitting the algorithm’s run into two phases. Let $\alpha_{\theta}$ be the last (largest) time $t$ such that $p_{t}<\theta$ (If this never happens, then we take $\alpha_{\theta}=0$). By the end of the first phase, i.e. at time $\alpha_{\theta}$, there are no pending transactions whose value per unit is at least $\theta$ (since these should have been scheduled before the transaction whose value per unit is $p_{\alpha_{\theta}}$). Thus, during the first phase, the total size of transactions with value at least $\theta$ that ALG schedules is at least as high as the total scheduled by OPT, that is, $Q_{[1,\alpha_{\theta}]}(ALG,\theta)\geq Q_{[1,\alpha_{\theta}]}(OPT,\theta)$. If $\alpha_{\theta}\geq T$, the claim follows. Otherwise, $\alpha_{\theta}<T$. Now we look at the second phase, i.e. during times steps $\alpha_{\theta}+1,\ldots,T,T+1$. For any time $t$ in this time range it holds that $p_{t}\geq\theta$, and thus every transaction scheduled by the algorithm in that range has value per unit of at least $\theta$. Since, as mentioned above, in every $Z$ consecutive blocks the algorithm schedules transactions of size at least $(Z-1)\cdot B$, we have that during the second stage (including time $T+1$) our algorithm scheduled transactions of size at least $((T+1)-\alpha_{\theta}-1)\cdot B$, while $OPT$ can schedule transactions of size at most $(T-\alpha_{\theta})\cdot B$ during the range $\alpha_{\theta}+1,\ldots,T-1,T$ (i.e without the extra step at time $(T+1)$). It follows that in the second phase, the total size of transactions with value at least $\theta$ that ALG schedules (with the extra time step) is at least as high as the total scheduled by OPT, that is, $Q_{[\alpha_{\theta}+1,T+1]}(ALG,\theta)\geq(T-\alpha_{\theta})\cdot B\geq Q_{[\alpha_{\theta}+1,T]}(OPT,\theta)$. ∎
The proposition can be directly deduced from this lemma using Lemma 2.6 that states that the social welfare achieved by $ALG$ for time $1$ till $t+1$ can be expressed as $\int_{0}^{\infty}Q_{[1,t+1]}(ALG,\theta)d\theta$, and similarly, the social welfare achieved by $OPT$ for time $1$ till $t$ can be expressed as $\int_{0}^{\infty}Q_{[1,t]}(OPT,\theta)d\theta$. ∎
@@ 3.2 Our Main Result
Our main result is that EIP-1559 algorithms have the attractive property that for patient bidders (with some bounds on the per-unit valuations), for any time horizon, they give close to optimal welfare with respect to the benchmark of all schedules that are restricted to maximum block size $B$ and are running for slightly fewer steps.
This may be viewed as surprising since the algorithm is both restricted to online decisions, and furthermore, is restricted to be price-based, and yet competes with all offline clairvoyant schedules. On the other hand, it is not limited by a worst-case block size limit $B$, but rather block size limit $B$ holds only on the average with some slackness $\Delta$ (as shown in Proposition 2.11). Additionally, the algorithm gets an extension of additional $\Gamma$ time steps, for some appropriately chosen $\Gamma$ (that does not grow with the time $T$).
@@
1 is a corollary of a more general and stronger result that we prove. First, instead of the benchmark being only integral allocations, we can take the benchmark to be all fractional allocations (in which a transaction with per-unit value $v_{i}$ that is allocated a fraction $x_{i}\in[0,1]$ contributes $v_{i}\cdot q_{i}\cdot x_{i}$ to the value of the schedule). Second, we present tighter results for the case that $q_{max}<B$, replacing the assumption that $c>2$ with the assumption that $c>1+q_{max}/B$. Third, instead of the benchmark having worst-case block size limit $B$, we allow it to only have average block size limit $B$ with slackness $\Delta^{\prime}$. To allow for this extra slackness we will allow the algorithm an additional $\Delta^{\prime}$ time steps (effectively increasing $\Gamma$ by $\Delta^{\prime}$). Finally, we will allow the initial price $p_{1}$ to be arbitrary. When it is larger than $v_{max}$ it will imply that $\Gamma$ grows as $\ln\frac{p_{1}}{p_{min}}$.
@@
We remark that if $p_{min}>v_{min}\cdot e^{-\eta}$ then the result still holds, but with respect to schedules that are only allowed to schedule transactions with value per unit at least $p_{min}\cdot e^{\eta}$. That is, there is an additional additive loss in welfare that is bounded by the welfare of any schedule of transactions of value per unit less than $p_{min}\cdot e^{\eta}$.
The cornerstone of the proof of the main result is the next lemma which shows that for any $\theta$, the EIP-1559 algorithm schedules large enough size of transactions of per-unit value $\theta$, as long as it gets a long enough extension. The theorem will be easily deduced from this bound using Lemma 2.6. For the next lemma, recall that $Q_{t}(S,\theta)$ denotes the total size of transactions with per-unit value of at least $\theta$ that were scheduled by $S$ at time $t$.
@@
The theorem easily follows from the main lemma by integrating Eq. 1 over $\theta$, and then applying Lemma 2.6. By integrating Eq. 1 we conclude that
$\int_{0}^{\infty}Q_{[1,T]}(S,\theta)d\theta\leq\int_{0}^{\infty}Q_{[1,T+\Gamma]}(E,\theta\cdot e^{-\eta})d\theta$ (2)
By Lemma 2.6 the LHS of Eq. 2 is exactly $SW(S,[1,T])$. To evaluate the RHS of Eq. 2, we make a linear change of variable (with our notation reusing $\theta$) $\int_{0}^{\infty}Q_{[1,T+\Gamma]}(E,\theta\cdot e^{-\eta})d\theta=e^{\eta}\int_{0}^{\infty}Q_{[1,T+\Gamma]}(E,\theta)d\theta$, and we use Lemma 2.6 again to conclude that the RHS is exactly $e^{\eta}\cdot SW(E,[1,T+\Gamma])$, as needed.
@@ 3.3 Proof of The Main Lemma
Before giving a formal proof of Lemma 3.3, let us give the intuition drawing parallels with the online algorithm above. Similarly to the previous analysis we will look at the last point in time $\alpha_{\theta}$ (to be simply called $\alpha$ below) where the price was less than $\theta$ and the quantity scheduled is at most $Q_{t}\leq c\cdot B-q_{max}$, meaning that every transaction with per-unit value at least the price was scheduled. As previously, up to that point, the EIP-1559 algorithm scheduled everything with per-unit value of at least $\theta$, so it competes with any schedule $S$ (and also with a fractional one). From this point on there are always enough pending transactions that are willing to pay at least $\theta$ per-unit (quantity larger than $c\cdot B-q_{max}$), so EIP-1559 prices will increase rapidly until they are at least $\theta^{\prime}=\theta\cdot e^{-\eta}$ (at a time to be called $\beta$ below), from which point they can never decrease much below $\theta^{\prime}$ (prices will always be at least $\theta^{\prime}=\theta\cdot e^{-\eta}$, the minimum price when price decreases from $\theta$ in a single step). So after this point in time, the EIP-1559 algorithm fills its quota with transactions of value per unit of at least $\theta\cdot e^{-\eta}$, and since its average block size limit is the same as that of $S$, it competes with it successfully given enough scheduling time. To compensate for the time it took for the price to rise up to at least $\theta^{\prime}$ (the time between $\alpha$ and $\beta$), we allow the algorithm an extended time of extra $\Gamma$ steps to schedule transactions. Additionally, we allow an extra $\Delta^{\prime}$ steps to the algorithm, in order for it to compete also with schedules that have average block size limitations with slackness $\Delta^{\prime}$. We suffer a loss of a $1-e^{-\eta}$ factor due to the possible single price decrease below $\theta$.
@@
To simplify notation we use $S_{t}(\theta)=Q_{t}(S,\theta)$ and $E_{t}(\theta)=Q_{t}(E,\theta)$. Let $Z_{t}(\theta)$ denote the total size of all transactions that arrived by time $t$ with value at least $\theta$, i.e. the size of transaction $i$ is counted towards $Z_{t}(\theta)$ if $v_{i}\geq\theta$ and $t_{i}\leq t$. It holds that
@@
Thus, in the rest of the proof we assume that demand at price $\theta$ was not necessarily met at any time $t\in[\alpha+1,T+\Gamma]$, and move to lower bound the total capacity scheduled by the EIP-1559 algorithm between time $\alpha+1$ and $T+\Gamma$. As demand at price $\theta$ was not necessarily met at any time $t\in[\alpha+1,T+\Gamma]$, then for $t$ in this range of times, whenever $p_{t}\leq\theta$ then $Q_{t}>c^{\prime}\cdot B\geq B$. We conclude that if $p_{t}\leq\theta$ then $p_{t+1}=p_{t}\cdot e^{\eta\cdot(Q_{t}-B)/B}>p_{t}\cdot e^{\eta\cdot(c^{\prime}-1)}$. Thus, if $p_{\alpha}<\theta$ then price will start rising till at some time $t\geq\alpha+1$ it will hold that $p_{t}\geq\theta\geq\theta^{\prime}$.
@@
Now, we use $\beta\in[\alpha+1,T+\Gamma]$ to denote the first time the price was at least $\theta^{\prime}$ (the existence of such a time is implied by the arguments below since $\Gamma$ is large enough). So, if $\beta>{\color[rgb]{1,0,0}\alpha+}1$ then it holds that $p_{\beta}\geq\theta^{\prime}$ and $p_{\beta-1}<\theta^{\prime}$ (note that for any time $t\in[\beta,T+\Gamma]$ it holds that $p_{t}\geq\theta^{\prime}$).
@@
Observe that the leftmost inequality is trivial as $\beta\geq\alpha+1$. Let us now consider the rightmost inequality. It holds due to the average block size limitation constraint of $B$ with slackness $\Delta^{\prime}$, which states that $\sum_{t=\alpha+1}^{T}S_{t}(\theta)\leq(T-\alpha+\Delta^{\prime})\cdot B$. The equality $\sum_{t=\beta}^{T+\Gamma}E_{t}(\theta^{\prime})=\sum_{t=\beta}^{T+\Gamma}Q_{t}$ holds as for any time $t\in[\beta,T+\Gamma]$ it holds that $p_{t}\geq\theta^{\prime}$, and thus all capacity is used to schedule transactions with value per unit at least $\theta^{\prime}$.
@@
Let us first get two special cases out of the way. The first special case is the case that $\beta=1$ and $\alpha=0$. In this case $p_{\beta}=p_{1}$ and $p_{{T+\Gamma+1}}\geq p_{min}$. So
@@
where the left inequality holds by the definition of $\Gamma$, and the claim follows. Thus, in the rest of the proof we assume that it is not the case that $\beta=\alpha+1=1$.
The second special case is $\beta=\alpha+1>1$. In this case
$p_{\beta}=p_{\alpha+1}=p_{\alpha}\cdot e^{\eta\cdot(Q_{\alpha}-B)/B}\leq\theta\cdot e^{\eta\cdot(c^{\prime}-1)}$
where the inequality holds since $p_{\alpha}\leq\theta$ and $Q_{\alpha}\leq c^{\prime}B$. As once the price gets to $\theta^{\prime}$ it never drops below it, we have $p_{T+\Gamma+1}\geq\theta^{\prime}$. We conclude that
$\frac{p_{\beta}}{p_{T+\Gamma+1}}\leq\frac{\theta\cdot e^{\eta\cdot(c^{\prime}-1)}}{\theta^{\prime}}=e^{\eta\cdot c^{\prime}}$
and therefore
$\frac{1}{\eta}\ln\left(\frac{p_{\beta}}{p_{{T+\Gamma+1}}}\right)\leq\frac{1}{\eta}\ln(e^{\eta\cdot c^{\prime}})=c^{\prime}$
As $\beta-(\alpha+1)=0$, we get
$\frac{1}{\eta}\ln\left(\frac{p_{\beta}}{p_{{T+\Gamma+1}}}\right)+\Delta^{\prime}+\beta-(\alpha+1)\leq c^{\prime}+\Delta^{\prime}.$ (4)
By the assumption about $\Gamma$,
$\Gamma\geq\frac{1}{\eta(c^{\prime}-1)}\cdot\ln\left(\frac{v_{max}}{p_{min}}\right)+c-1+\frac{c-2}{c^{\prime}-1}+\Delta^{\prime}.$
Since $v_{max}\geq v_{min}\geq e^{\eta}p_{min}$, we have
$\frac{1}{\eta(c^{\prime}-1)}\ln\left(\frac{v_{max}}{p_{min}}\right)\geq\frac{1}{c^{\prime}-1}.$
Therefore
$\Gamma\geq\Delta^{\prime}+c-1+\frac{1}{c^{\prime}-1}+\frac{c-2}{c^{\prime}-1}=\Delta^{\prime}+c-1+\frac{c-1}{c^{\prime}-1}.$
As $c^{\prime}\leq c$, we have $c^{\prime}-1\leq c-1$, and hence
$\frac{c-1}{c^{\prime}-1}\geq 1.$
So
$\Gamma\geq\Delta^{\prime}+c\geq\Delta^{\prime}+c^{\prime},$
and the claim follows by Eq. 4.
We now analyze the general case where $\beta>\alpha+1\geq 1$. Recall that $\beta\in[\alpha+1,T+\Gamma]$ is the first time such that $p_{\beta}\geq\theta^{\prime}$. Thus, as $\beta>\alpha+1$ it holds that $p_{\beta-1}<\theta^{\prime}$ (note that we have shown that for any time $t\in[\beta,T+\Gamma]$ it holds that $p_{t}\geq\theta^{\prime}$), and as $p_{\alpha+1}\geq p_{min}$ we get
@@
As $\eta>0$ and $c>c^{\prime}>1$ (so c2>1)c-2>-1) it holds that $v_{max}\cdot e^{\eta\cdot(c-2)}>v_{max}\cdot e^{-\eta}\geq\theta\cdot e^{-\eta}\geq\theta^{\prime}$, by Eq. 5